# Unveiling the Potential of iMarkers: Invisible Fiducial Markers for Advanced Robotics

### v4

@@ 5.4 Recognition Speed
Considering the necessity of fast recognition speeds in demonstrating the practicality of fiducial markers for real-time applications, this subsection provides a detailed profiling analysis of iMarker detection and recognition. Table 3 summarizes cumulative processing times for iMarker data decoding. Evaluations were conducted using various sensor configurations and implemented software introduced in subsection 5.1. The algorithms developed for all sensor variants share five core modules: acquisition,” which captures video frames; preprocessing,” which applies image enhancement techniques; processing,” which refines frames to emphasize marker patterns; post-processing,” which enhances the final binary images; and recognition to detect and interpret marker patterns. To obtain the measurements, $7\times 7~\mathrm{cm^{2}}$ iMarkers were placed approximately $50~\mathrm{cm}$ from the lens, directly facing the sensor under standard lighting conditions with evenly diffused artificial illumination. For the UV-range iMarker, a DARKBEAM A300 365nm365~nm flashlight was additionally used to enhance the iMarker visibility.
According to Table 3, the overall iMarker recognition process, from video frame acquisition to inner pattern recognition steps, is completed within milliseconds. To compare, some methods discussed in Table 1 require several seconds for detection, highlighting the efficiency of the iMarker system. Among the various configurations, UV-range iMarker recognition (Algorithm 4) is the fastest, as the input image is already in grayscale, requiring only a simple “thresholding” operation to reveal the marker pattern for recognition. In dual-vision setups (Algorithm 1), acquisition and processing account for the subsequent significant portions of the processing time due to the need to receive synchronized frames from two cameras and perform image alignment for further processing, respectively. The post-processing stage is executed more efficiently, as the substantial filtering process retains only the differing image regions, i.e., the CSR areas. On the other hand, single-vision setups operate faster in acquisition due to capturing frames from a single camera. However, they allocate more time to post-processing, as the captured images contain higher noise levels and the need to refine iMarker regions effectively.
6 Discussion
@@ 6.1 Comparative Insights
In Fig. 10-a, three primary performance aspects of the iMarker variants have been analyzed. Regarding “invisibility,” UV-range iMarkers remain entirely imperceptible to the eye, while IR-range iMarkers may exhibit slight visibility due to scattering under extreme lighting conditions. Visible-range iMarkers, on the other hand, are designed to blend into patterned backgrounds; however, upon close inspection, they remain minimally visible to the human eye. When “readability” is studied, visible-range iMarkers have better compatibility with standard cameras and do not require additional illumination. However, IR- and UV-range iMarkers require dedicated illumination sources and particular sensors, with varied detection performance based on environmental conditions. Finally, and in terms of “detection equipment cost” (camera, flashlight, etc.), visible-range iMarkers are the most economical, as they can be detected using standard cameras. UV equipment (camera and illumination sources) is generally the most expensive among the three, and IR-range iMarker detection equipment cost lies in between.
@@
Considering the sensor-related factors shown in Fig. 10-b, the performance scores can be studied in five aspects. In “design complexity,” the static single-vision setup is the most straightforward, requiring only a static polarizer attached to the lens. The dynamic single-vision setup has moderate complexity by incorporating a function generator to control polarization. Dual-vision configurations represent the most complex designs, with the cube beamsplitter offering more reliable light division than the plate one, which requires precise alignment for optimal beam splitting. The design complexity directly influences the “sensor cost.” The static single-vision setup offers the most economical solution, requiring only a polarizer, while the dynamic single-vision configuration incurs additional costs due to its polarization control components. Dual-vision setups are generally more expensive. They require two cameras, a beamsplitter (either cube or plate), and a structural mounting cage to ensure precise alignment and stability.
@@
Regarding the “field of view,” single-vision setups provide a broader theoretical FoV (e.g., $87^{\circ}\times 58^{\circ}$ in the RealSense™ D435, shown in Fig. 6-c). However, in dual-vision setups, the beamsplitter’s physical dimensions and the supporting cage partially obstruct the cameras’ FoV, reducing the effective viewing area. For instance, the configuration shown in Fig. 6-c achieves an effective FoV of $57^{\circ}\times 54.5^{\circ}$. Another essential factor is “weight”; the static single-vision setup is the lightest, whereas the dynamic single-vision setup adds approximately 30gr30~gr due to the function generator and dynamic polarizer. Thus, single-vision setups are more feasible for minimizing payloads, such as in drones, as dual-vision configurations are heavier. The combined weight of dual-vision setup components typically ranges from $150$ to 350gr350~gr, with the plate beamsplitter potentially reducing the overall sensor weight. Finally, considering the noise levels in the processed output, dual-vision setups produce clearer results, isolating only the CSR regions of the iMarkers. In contrast, the dynamic single-vision setup may introduce additional noise due to motion and illumination fluctuations during “temporal” subtraction. The static single-vision setup retains the highest noise pixels (i.e., pixels that do not contribute to iMarker detection), including the filtered areas with the detected iMarkers.
@@
When choosing among various sensor configurations, a selection parameter is the trade-off between “detection robustness” and “clutter rejection,” as each configuration’s detection pipeline can be tailored to optimize performance for specific application scenarios. Dual-vision setups rely on spatial frame subtraction followed by thresholding, producing cleaner outputs by isolating CSR-coated iMarker regions while effectively suppressing non-marker backgrounds thanks to the differential visibility of CSRs between the two sensor views. However, they are more sensitive to artifacts caused by sensor noise, motion, illumination variations, and frame misalignment. In contrast, single-vision setups may retain more background clutter (noise pixels) alongside detected iMarkers. Accordingly, enhancing robustness and output quality are tackled differently considering the employed sensor variant. While strategies such as adaptive thresholding, active frame registration, and noise modeling can better aid dual-vision setup algorithms, single-vision systems can benefit from lightweight morphological filtering to enhance detection quality without compromising system simplicity.
6.2 Applicability
@@
Considering the unique characteristics of iMarkers, they serve as potential alternatives in scenarios reliant on printed markers. In this regard, the obscure nature of iMarkers makes them appropriate for surroundings with unwanted visual clutter, such as museums, hospitals, offices, etc. By safely replacing printed markers, iMarkers can preserve unobtrusiveness and aesthetic integrity while delivering full functionality for robotics and AR systems. It should be noted that various AR applications, particularly those deployable on handheld mobile devices, rely exclusively on RGB cameras and frequently utilize fiducial landmarks to support interactive functionalities like user engagement. This aligns well with the iMarker detection procedure proposed in this paper; specifically the single-vision setup which requires minimal hardware modifications to enable compatibility with standard AR-capable vision sensors.
@@
Additionally, as the iMarker detection pipeline relies on simple computer vision techniques (e.g., thresholding and frame subtraction), it introduces minimal computational overhead and operates efficiently even on a CPU-based system, without requiring hardware acceleration. Thus, the core innovation of the proposed system lies in tightly coupling hardware and software, where the sensor hardware captures frames with filtered views on CSR-coated regions of iMarkers and the software extracts and reveals these regions before passing the result to standard fiducial marker libraries (e.g., ArUco or AprilTag).
Fig. 11 depicts the real-world benchmarking of iMarkers in Visual Simultaneous Localization and Mapping (VSLAM), showcasing the use of a visible-range iMarker to facilitate indoor map reconstruction enriched with semantic information [20]. iMarkers can also be suitable for autonomous robots functioning in retail stores and smart warehousing, where numerous markers are needed to augment objects. Another potential use case includes object grasping in cluttered environments, where iMarkers can supply accurate localization cues to enhance robotic manipulation accuracy. iMarkers can also be strategically employed to seamlessly augment difficult-to-detect objects, like transparent surfaces (e.g., windows and glass block walls), where the transparency poses considerable challenges to computer vision and robot perception systems. Likewise, reflective surfaces, such as mirrors or chrome/metallic objects, often cause distortions or reflections, confusing perception algorithms and negatively impacting object recognition or localization accuracy. In these scenarios, iMarkers can augment such objects by providing reliable information while preserving visual appeal. This allows robots and AR devices to effectively see and interact with visually challenging entities, ensuring reliable operation and versatility.
Figure 11: Benchmarking iMarkers for map reconstruction in alliance with printed markers introduced in [20].
@@ 6.3 Limitations and Fabrication Considerations
Given the ongoing nature of this research and the absence of a rapid iMarker fabrication process so far, the experiments remain constrained under varying illumination, weather conditions (for outdoor applications), and environmental factors (e.g., dirt). In terms of environmental robustness, iMarkers are expected to exhibit similar vulnerability to dirt and occlusion as conventional printed markers, since both rely on visual input for marker detection and pose estimation. In other words, the marker recognition pipeline (ArUco in this work) processes either standard RGB frames of normal cameras or the binary outputs of the iMarker detector algorithm, without introducing additional sensitivity. Regarding sunlight and outdoor use, UV-range iMarkers may experience reduced contrast under ambient sunlight, potentially leading to false positives due to uncontrolled UV reflections. Accordingly, the reflectivity of UV-range CSRs needs to be enhanced to ensure that all UV-range iMarkers are reliably detected when illuminated by the light source. Similarly, IR-range iMarkers could be influenced by environmental IR sources. For these cases in real-world deployments, additional filtering strategies or controlled illumination may be required to ensure reliable iMarker detection.
The scalability of iMarker fabrication for mass deployment is still an active investigation topic. Currently, production involves several manual steps, including microfluidic droplet generation, annealing, polymerization, and sequential washing, which reduce throughput and introduce variability in marker quality. To overcome this, we are developing an integrated, automated production platform that combines all stages into a continuous process with real-time computer monitoring. This system is expected to streamline manufacturing, improve consistency, and enable large-scale high-throughput fabrication suitable for real-world robotics applications.
@@ 7 Conclusions
This paper introduced a novel generation of fiducial markers, termed iMarkers, designed explicitly for real-world robotics applications. iMarkers are designed to be visually indistinguishable by the human eye, ensuring minimal environmental disruption while remaining fully detectable and recognizable by robots equipped with adequate sensors and adapted detection algorithms. The paper also presents various sensor designs and implemented algorithms for detecting iMarkers placed in the environment and exploiting them for different use cases. Experimental results demonstrate the practicality and potential of these markers as versatile landmarks for data embedding in various real-world scenarios.
Future works include optimizing the iMarker fabrication process to ensure scalable production and enhancing the illumination equipment, sensor designs, and algorithm implementations to guarantee robustness across various scenarios.
@@ Acknowledgments
The authors would like to thank Marco Giberna and Tadej Emersic for their insightful assistance during experiments conducted in this work.

### v5

@@ 5.5 Recognition Speed
Given the need for fast recognition to demonstrate the practicality of fiducial markers for real-time applications (in our context, with an average processing rate of $25\pm 5\mathrm{fps}$), this subsection presents a detailed profiling analysis of iMarker detection and recognition. Table 4 summarizes cumulative processing times for iMarker data decoding. Evaluations were conducted using various sensor configurations and implemented software introduced in subsection 5.1. The algorithms developed for all sensor variants share five core modules: acquisition,” which captures video frames; preprocessing,” which applies image enhancement techniques; processing,” which refines frames to emphasize marker patterns; post-processing,” which enhances the final binary images; and recognition to detect and interpret marker patterns. To obtain the measurements, $7~\mathrm{cm}\times 7~\mathrm{cm}$ iMarkers were placed approximately $50~\mathrm{cm}$ from the lens, directly facing the sensor under standard lighting conditions with evenly diffused artificial illumination. For the UV-range iMarker, a DARKBEAM A300 $365~\mathrm{nm}$ flashlight was additionally used to enhance the iMarker visibility.
According to Table 4, the overall iMarker recognition process, from video frame acquisition to inner pattern recognition steps, is completed within milliseconds and adds only $\approx 100~ms$ overhead to the ArUco library. The reported results are obtained using standard CPU-based processing and without the aid of GPU or other hardware accelerators. For comparison, some methods discussed in Table 1 require several seconds for detection, underscoring the efficiency of the iMarker system. Among the various configurations, UV-range iMarker recognition (Algorithm 4) is the fastest, as the input image is already in grayscale, requiring only a simple “thresholding” operation to reveal the marker pattern for recognition. In dual-vision setups (Algorithm 1), acquisition and processing account for the subsequent significant portions of the processing time due to the need to receive synchronized frames from two cameras and perform image alignment for further processing, respectively. The post-processing stage is executed more efficiently because the substantial filtering process retains only the regions that differ, i.e., the CSR areas. On the other hand, single-vision setups are faster during acquisition because they capture frames from a single camera. However, they allocate more time to post-processing, as the captured images contain higher noise levels and require effective refinement of iMarker regions.
6 Discussions
@@ 6.1 Comparative Insights
In Fig. 10a, three primary performance aspects of the iMarker variants have been analyzed. Regarding “invisibility,” UV-range iMarkers remain entirely imperceptible to the eye, while IR-range iMarkers may exhibit slight visibility due to scattering under extreme lighting conditions. Visible-range iMarkers, on the other hand, are designed to blend into patterned backgrounds; however, upon close inspection, they remain minimally visible to the human eye. When “readability” is studied, visible-range iMarkers have better compatibility with standard cameras and do not require additional illumination. However, IR- and UV-range iMarkers require dedicated illumination sources and specific sensors, and their detection performance varies with environmental conditions. Finally, and in terms of “detection equipment cost” (camera, flashlight, etc.), visible-range iMarkers are the most economical, as they can be detected using standard cameras. UV equipment (camera and illumination sources) is generally the most expensive of the three, and IR-range iMarker detection equipment costs lie between them.
@@
Considering the sensor-related factors shown in Fig. 10b, the performance scores can be studied in five aspects. In “design complexity,” the static single-vision setup is the most straightforward, requiring only a static polarizer attached to the lens. The dynamic single-vision setup has moderate complexity, incorporating a function generator to control polarization. Dual-vision configurations are more complex, with the cube beamsplitter offering more reliable light division than the plate beamsplitter, which requires precise alignment for optimal splitting. The design complexity directly influences the “sensor cost.” The static single-vision setup is the most economical, requiring only a polarizer, whereas the dynamic single-vision configuration incurs additional costs due to its polarization-control components. Dual-vision setups are generally more expensive: they require two cameras, a beamsplitter (either cube or plate), and a structural mounting cage to ensure precise alignment and stability.
@@
Regarding the “field of view,” single-vision setups provide a broader theoretical FoV (e.g., $87^{\circ}\times 58^{\circ}$ in the RealSense™ D435, shown in Fig. 6c). However, in dual-vision setups, the beamsplitter’s physical dimensions and the supporting cage partially obstruct the cameras’ FoV, reducing the effective viewing area. For instance, the configuration shown in Fig. 6-c achieves an effective FoV of $57^{\circ}\times 54.5^{\circ}$. Another essential factor is “weight”; the static single-vision setup is the lightest, whereas the dynamic single-vision setup adds approximately $30~\mathrm{gr}$ due to the function generator and dynamic polarizer. Thus, single-vision setups are more feasible for minimizing payload weight, e.g., in drones, as dual-vision configurations are heavier. The combined weight of dual-vision setup components typically ranges from $150$ to $350~\mathrm{gr}$, with the plate beamsplitter potentially reducing the overall sensor weight.
@@
When choosing among sensor configurations, a key selection criterion is the trade-off between “detection robustness” and “clutter (noise) rejection,” as each configuration’s detection pipeline can be tailored to optimize performance for specific application scenarios. Dual-vision setups rely on spatial frame subtraction followed by thresholding, producing cleaner outputs by isolating CSR-coated iMarker regions while effectively suppressing non-marker backgrounds thanks to the differential visibility of CSRs between the two sensor views. However, they are more sensitive to artifacts caused by sensor noise, motion, illumination variations, and frame misalignment. In contrast, single-vision setups may retain more background clutter (noise pixels) alongside detected iMarkers. Accordingly, enhancing robustness and output quality is addressed differently depending on the sensor variant. While strategies such as adaptive thresholding and active frame registration can better aid dual-vision setup algorithms, single-vision systems can benefit from lightweight morphological filtering to enhance detection quality without compromising system simplicity.
6.2 Potential Applications
@@
Considering the unique characteristics of iMarkers, they serve as potential alternatives in scenarios reliant on printed markers. In this regard, the obscure nature of iMarkers makes them appropriate for surroundings with unwanted visual clutter, such as museums, hospitals, offices, etc., and for scenarios involving transparent or reflective structures (e.g., glass walls and partitions), where placing printed fiducial markers is impractical or unacceptable due to their high visibility and disruption of visual transparency. By safely replacing printed markers, iMarkers can preserve unobtrusiveness and aesthetic integrity while delivering full functionality for robotics and AR systems. Note that various AR applications, particularly those deployed on handheld mobile devices, rely exclusively on RGB cameras and often use fiducial landmarks to support interactive features such as user engagement. This aligns well with the iMarker detection procedure proposed in this paper; specifically, the single-vision setup, which requires minimal hardware modifications to enable compatibility with standard AR-capable vision sensors.
@@
Additionally, as the iMarker detection pipeline relies on simple computer vision techniques (e.g., thresholding and frame subtraction), it introduces minimal computational overhead and operates efficiently even on a CPU-based system, without requiring hardware acceleration. Thus, the core innovation of the proposed system lies in tightly coupling hardware and software: the sensor hardware captures frames with filtered CSR-coated iMarker regions, and the software extracts and reveals these regions before passing the results to standard fiducial marker libraries (e.g., ArUco or AprilTag).
Fig. 11 depicts two representative real-world robotics applications in which iMarkers play a crucial role. In Fig. 11(a), iMarkers are benchmarked within a Visual Simultaneous Localization and Mapping (VSLAM) framework, showcasing the use of visible-range landmarks to support indoor map reconstruction enriched with semantic information [17]. Accordingly, iMarkers are applied to office door frames to unobtrusively encode semantic cues related to structural entities, thereby enabling semantic mapping while remaining visually camouflaged within the environment. In another setting, Fig. 11(b) shows a UAV-based robot-landing scenario introduced in Section 5.4, where conventional printed markers become unreliable or undetectable under low-illumination conditions, while IR-range iMarkers remain robustly detectable, enabling reliable localization and safe landing despite severely degraded visible-light cues.
((a)) iMarkers integrated into a visual SLAM pipeline to support camera localization and semantic annotation of indoor environments [17].
((b)) iMarkers deployed as low-illumination landing targets, enabling reliable pose estimation and precise autonomous landing of aerial robots.
Figure 11: Real-world robotics applications enabled by iMarkers.
@@ 6.3 Limitations and Fabrication Considerations
Given the ongoing nature of this research and the absence of a rapid iMarker fabrication process to date, the experiments remain constrained by varying illumination conditions, weather (for outdoor applications), and environmental factors (e.g., dirt). In terms of environmental robustness, iMarkers are expected to be susceptible to dirt and occlusion, as with conventional printed markers, since both rely on visual input for detection and pose estimation. In other words, the marker recognition pipeline (ArUco in this work) processes either RGB frames from standard cameras or the binary outputs of the iMarker detector, without introducing additional sensitivity. Regarding sunlight and outdoor use, UV-range iMarkers may experience reduced contrast under ambient sunlight, potentially leading to false positives due to uncontrolled UV reflections. Accordingly, the reflectivity of UV-range CSRs needs to be enhanced to ensure that all UV-range iMarkers are reliably detected when illuminated by the light source. Similarly, IR-range iMarkers could be influenced by environmental IR sources. In real-world deployments, additional filtering strategies or controlled illumination may be required to ensure reliable iMarker detection.
The scalability of iMarker fabrication for mass deployment remains an active area of investigation. Currently, production involves several manual steps, including microfluidic droplet generation, annealing, polymerization, and sequential washing, which reduce throughput and introduce variability in marker quality. To overcome this, we are developing an automated production platform that combines all stages into a continuous process with real-time computer monitoring. This system is expected to streamline manufacturing, improve consistency, and enable large-scale high-throughput fabrication suitable for real-world robotics applications.
@@ 7 Conclusions
This paper introduced a novel generation of fiducial markers, termed iMarkers, designed explicitly for real-world robotics applications. iMarkers are designed to be visually indistinguishable to the human eye, ensuring minimal environmental disruption while remaining fully detectable and recognizable by robots equipped with appropriate sensors and detection algorithms. The paper also presents various sensor designs and implemented algorithms for detecting iMarkers placed in the environment and exploiting them for different use cases. Experimental results demonstrate the practicality and potential of these markers as versatile landmarks for data embedding in various real-world scenarios.
Future work includes optimizing the iMarker fabrication process to ensure scalable production and enhancing illumination equipment, sensor designs, and algorithm implementations to ensure robustness across various scenarios.
@@ Acknowledgments
The authors would like to thank Marco Giberna, Dr. Tadej Emersic, Dr. Xu Ma, and Aabhash Bhandari for their assistance in conducting experiments.